OpenStudy (thedj4jc):

Solve for m. Sqrt m-7=n+3 m = n2 + 6n + 2 m = n2 + 6n - 2 m = n2 + 6n + 16

OpenStudy (thedj4jc):

\[\sqrt{m-7}=n+3\]

OpenStudy (butterflydreamer):

first square both sides :) This will get rid of the sqrt sign.

OpenStudy (butterflydreamer):